If a Triangle's height is 4 inches less than two times its base, and has an area of 120 square inches. What is the height of the triangle?
Answer by nerdybill(7384)

If a Triangle's height is 4 inches less than two times its base, and has an area of 120 square inches. What is the height of the triangle?
.
Let b = base
then
2b-4 = height
.
area = (1/2)bh
120 = (1/2)b(2b-4)
240 = b(2b-4)
240 = 2b^2-4b
120 = b^2-2b
0 = b^2-2b-120
0 = (b-12)(b+10)
b = 12 inches (base)
.
height:
2b-4 = 2(12)-4 = 24-4 = 20 inches